# Quantitative Propagation of Chaos for 2D Viscous Vortex Model with General Circulations on the Whole Space

### v1

@@ 2.5. Solving the New ODE Hierarchy
This has given us a key idea that the iterated integral itself solves another much simpler ODE hierarchy, which enables us this time to use the Grönwall’s inequality and direct induction to derive the explicit expressions of the complicated integrals. In order to cancel the first term in the right-hand side, we denote that
@@
which solves by simple calculations that
@@
$\displaystyle A_{k}^{l}(t_{k})=\Big{(}k(k+1)...l\Big{)}\int_{0}^{t_{k}}...\int_{0}^{t_{l}}$ $\displaystyle\exp\Big{(}-\sum_{j=k}^{l}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}\Big{)}$
$\displaystyle f(t_{k+1})...f(t_{l+1})\,\mathrm{d}t_{l+1}...\,\mathrm{d}t_{k+1}.$
@@
$\displaystyle A_{k}^{l}(t_{k})=$ $\displaystyle\,\Big{(}k(k+1)...(l-1)\Big{)}\int_{0}^{t_{k}}...\int_{0}^{t_{l-1}}\exp\Big{(}-\sum_{j=k}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}$
$\displaystyle\,\qquad\qquad\qquad\qquad\qquad\qquad\Big{(}\exp(l\varphi(t_{l}))-1\Big{)}f(t_{k+1})...f(t_{l})\,\mathrm{d}t_{l}...\,\mathrm{d}t_{k+1}$
@@
Using the simple combinatory formula
@@
$\displaystyle\,=e^{-k\varphi(t)}\frac{(-1)^{l-k}}{(l-k)!}\Big{(}k(k+1)...l\Big{)}\int_{0}^{x}(y-x)^{l-k}(1-y)^{-l-1}\,\mathrm{d}y\Big{|}_{x=1-e^{-\varphi(t)}}.$
@@
where $Y$ is some random variable with the Beta distribution $B(k,l-k+1)$. By the sub-Gaussian tail estimate of the Beta distribution in [MA17, Theorem 2.1], which reads
@@ 3. Regularity Estimates
After presenting our proof of the main results, we are now turning to provide the key logarithmic growth estimates of the limit density, which are essentially applied in the previous discussions. In this section, we shall first collect some basic regularity estimates of the limit density $g$ as ingredients of our logarithmic estimates.
First we borrow some long-time asymptotic decay regularity estimates of the vorticity $\omega$ from Kato [Kat94], since $\omega$ solves the vorticity formulation of the 2D Navier-Stokes equation with an $L^{1}$ initial data, as pointed out before. This will give us enough knowledge about the coefficient of the drift term for the limit equation (1.7).
@@
Assume that $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}({\mathbb{R}}\times{\mathbb{R}}^{2})$ solves the nonlinear Fokker-Planck limit equation (1.7), then the vorticity $\omega(t,x)$ given by the formula
$\omega(t,x)=\int_{\mathbb{R}}mg(t,m,x)\,\mathrm{d}m$
satisfies the asymptotic decay estimates
$\|\nabla^{k}K\ast\omega(t,\cdot)\|_{\infty}\leq C(t\vee 1)^{-(1+k)/2}$
for any $k\geq 0$.
@@
The long time case is directly from [Kat94, Theorem II] and the short time case is proved by the parabolic maximum principle in [FW23, Lemma 2.2]. ∎
Next we shall apply the classical Carlen-Loss argument [CL96] to provide some Gaussian upper bounds and $L^{p}$ norm bounds with respect to the variable $x$ on the limit density, especially giving the explicit decay rate of those quantities in time $t$, which are optimal if one checks with the easiest heat kernel case. This result is due to the fact that the drift term $K\ast\omega$ is divergence-free and satisfies the long-time asymptotic decay rate as in the above lemma.
@@
Let $g$ solves the nonlinear Fokker-Planck limit equation (1.7). Assume on the initial data as in our main results that $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}({\mathbb{R}}\times{\mathbb{R}}^{2})$ satisfies the Gaussian upper bound
@@
Note that the diffusion coefficient $K\ast\omega$ in (1.7) is divergence-free, with
$\|K\ast\omega\|_{\infty}\leq Ct^{-1/2}$
by Lemma 3.1. Hence we can apply the optimal spatial decay result [CL96, Theorem 3] with $\beta=1/2$ for instance, say
$g(t,m,x)\leq C\int_{{\mathbb{R}}^{2}}\frac{1}{t}\exp\Big{(}-\frac{|x-y|^{2}}{8t}\Big{)}g_{0}(m,y)\,\mathrm{d}y.$
@@
$\displaystyle g(t,m,x)\leq\,$ $\displaystyle C\int_{{\mathbb{R}}^{2}}\frac{1}{t}\exp\Big{(}-\frac{|x-y|^{2}}{8t}-\frac{|y|^{2}}{C_{0}}\Big{)}\,\mathrm{d}y$
@@
The long-time $L^{p}$ norm control is obtained by the optimal smoothing result [CL96, Theorem 1], while the short-time $L^{p}$ norm control follows from the interpolation inequality and the fact that
$\|g(t,\cdot)\|_{1}\equiv 1,\quad\|g(t,\cdot)\|_{\infty}\leq\|g_{0}\|_{\infty}\leq C$
@@
Then we apply the classical parabolic maximum principle (see for instance the classical book by Evans [Eva22]) to provide some Gaussian lower bound on the limit density.
@@
Let $g$ solves the nonlinear Fokker-Planck limit equation (1.7). Assume on the initial data as in our main results that $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}({\mathbb{R}}\times{\mathbb{R}}^{2})$ satisfies the logarithmic gradient estimate
@@
Our idea is to construct an auxiliary function $\varphi(t,x)$ of some Gaussian type satisfying that
$\partial_{t}\varphi+(K\ast\omega)\cdot\nabla\varphi-\sigma\Delta\varphi\leq 0.$
Hence we conclude by the parabolic maximum principle that $\varphi(t,\cdot)-g(t,\cdot)\leq 0$ as long as $\varphi(0,\cdot)-g(0,\cdot)\leq 0$, thanks to that $\varphi-g$ satisfies the a priori exponential growth estimate. Assume that
$\varphi(t,x)=e^{\delta(t)}e^{\lambda(t)\frac{|x|^{2}}{2}}$
for some $\delta(t),\lambda(t)$ to be determined. Plug into the PDE, we have
$\displaystyle\,\partial_{t}\varphi+(K\ast\omega)\cdot\nabla\varphi-\sigma\Delta\varphi$
$\displaystyle$== $\displaystyle\,\Big{(}\delta^{\prime}+\lambda^{\prime}\frac{|x|^{2}}{2}+\lambda(K\ast\omega)\cdot x-2\sigma\lambda-\sigma\lambda^{2}|x|^{2}\Big{)}\varphi$
@@
Hence it suffices to let
$\lambda(t)=-\frac{C}{1+t}\text{ and }\delta(t)=-C-C\log(1+t).$
Now we check that $\varphi(0,x)=\exp\Big{(}-C-\frac{C}{2}|x|^{2}\Big{)}$, while our logarithmic gradient estimate assumption on the initial data together with the mean-value inequality yields that $g_{0}\geq\varphi(0,\cdot)$ for some large $C$. This will conclude our desired result. ∎
@@
Let $g$ solves the nonlinear Fokker-Planck limit equation (1.7). Assume on the initial data as in our main results that $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}({\mathbb{R}}\times{\mathbb{R}}^{2})$. Then we have for any $1\leq p\leq\infty$,
$\|\nabla g(t,m,\cdot)\|_{p}\leq Ce^{Ct},$
$\|\nabla^{2}g(t,m,\cdot)\|_{p}\leq C(1+t)e^{Ct}$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\nabla g\|_{\infty}\leq C\|K\ast\nabla\omega\|_{\infty}\|\nabla g\|_{\infty}\leq C\|\nabla g\|_{\infty},$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\nabla g\|_{1}\leq C\|K\ast\nabla\omega\|_{\infty}\|\nabla g\|_{1}\leq C\|\nabla g\|_{1}.$
@@
By the parabolic maximum principle, the evolution of $\|\nabla^{2}g\|_{\infty}$ is given by
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\nabla^{2}g\|_{\infty}\leq C\|K\ast\nabla\omega\|_{\infty}\|\nabla^{2}g\|_{\infty}+C\|K\ast\nabla^{2}\omega\|_{\infty}\|\nabla g\|_{\infty}\leq C\|\nabla^{2}g\|_{\infty}+Ce^{Ct},$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\nabla^{2}g\|_{1}\leq C\|K\ast\nabla\omega\|_{\infty}\|\nabla^{2}g\|_{1}+C\|K\ast\nabla^{2}\omega\|_{\infty}\|\nabla g\|_{1}\leq C\|\nabla^{2}g\|_{1}+Ce^{Ct}.$
@@
4. Logarithmic Estimates
In this section, we focus on the key logarithmic growth estimates stated before. As observed in many previous works, the major difficulty one will face when working on the whole space instead of the compact domain torus case, is the growth estimates for the derivatives of the logarithm of the limit density. Roughly speaking, when estimating the relative entropy dissipation, we need to set up some growth estimates for quantities like $\nabla\log\bar{\rho}$ and $\nabla^{2}\log\bar{\rho}$, where $\bar{\rho}$ represents generally the limit density. Such estimates are also crucial when considering the stability of the limit density in the sense of relative entropy, which goes beyond the range of particle approximation.
For the torus case, as in [JW18, GLBM24, SZ24], one may assume that the initial density has some positive lower bound $\lambda>0$. Hence by the parabolic maximum principle, the limit density keeps the lower bound for any time, giving the trivial estimate
@@ 4. Logarithmic Estimates
This assumption even leads to a time-uniform LSI for the limit density by the classical Bakry-Émery arguments, which plays the crucial role in [GLBM24] for proving the uniform-in-time propagation of chaos.
But for our whole space case, one can never find the positive lower bound for the density. Hence the authors introduced some ideas in [FW23] to set up the logarithmic gradient and Hessian estimates and started the program of extending the quantitative propagation of chaos results to the whole Euclidean space, where these results are obtained using the famous Li-Yau estimate [LY86] and Hamilton estimate [Ham93, HZ16], and the Bernstein method with Grigor’yan maximum principle [Gri06, Gri09, Li15]. A simple parabolic maximum principle method is available for the Landau-Maxwellian equation [CFG+24]. In this section we shall set up the logarithmic estimates for the equation (1.7), and we have simplified our method using only the Grigor’yan parabolic maximum principle, thanks to the Gaussian upper and lower bound results, compared to the previous work [FW23].
Proposition 4.1 (Logarithmic Estimates).
Assume that $g_{t}(m,x)\geq 0$ satisfies the nonlinear Fokker-Planck limit equation (1.7) with ×$\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}g_{t}(z)\,\mathrm{d}z=1$ for any $t$. Assume that $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}({\mathbb{R}}\times{\mathbb{R}}^{2})$ satisfies the logarithmic growth conditions
$|\nabla\log g_{0}(m,x)|^{2}\lesssim 1+|x|^{2}$
$|\nabla^{2}\log g_{0}(m,x)|\lesssim 1+|x|^{2}$
@@
$|\nabla\log g_{t}(m,x)|^{2}\lesssim\frac{1+\log(1+t)}{1+t}+\frac{|x|^{2}}{(1+t)^{2}},$ (4.1)
$|\nabla^{2}\log g_{t}(m,x)|\lesssim\frac{1+\log(1+t)}{1+t}+\frac{|x|^{2}}{(1+t)^{2}}.$ (4.2)
The proof of Proposition 4.1 is almost the same as in [FW23], namely applying the Bernstein’s method by creating some auxiliary functions and propagating them along the parabolic equation (1.7). Using the specific structure of the coefficients of the equation, we are able to apply the Grigor’yan parabolic maximum principle [Gri06, Gri09] to deduce the non-positivity of those auxiliary functions, leading to our desired estimates. For the sake of completeness, we present the proof again in the following subsections.

### v2

@@ 2.5. Solving the New ODE Hierarchy
This has given us a key idea that the iterated integral itself solves another much simpler ODE hierarchy, which enables us this time to use the Grönwall’s inequality and direct induction to derive the explicit expressions of the complicated integrals. In order to cancel the first term on the right-hand side, we denote that
@@
which solves by simple computations that
@@
$\displaystyle A_{k}^{l}(t_{k})=\Big{(}k(k+1)\cdots l\Big{)}\int_{0}^{t_{k}}\cdots\int_{0}^{t_{l}}$ $\displaystyle\,\exp\Big{(}-\sum_{j=k}^{l}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}\Big{)}$
$\displaystyle\,f(t_{k+1})\cdots f(t_{l+1})\,\mathrm{d}t_{l+1}\cdots\,\mathrm{d}t_{k+1}.$
@@
$\displaystyle A_{k}^{l}(t_{k})=$ $\displaystyle\,\Big{(}k(k+1)\cdots(l-1)\Big{)}\int_{0}^{t_{k}}\cdots\int_{0}^{t_{l-1}}\exp\Big{(}-\sum_{j=k}^{l-1}j\big{(}\varphi(t_{j})-\varphi(t_{j+1})\big{)}-l\varphi(t_{l})\Big{)}$
$\displaystyle\,\qquad\qquad\qquad\qquad\qquad\qquad\Big{(}\exp(l\varphi(t_{l}))-1\Big{)}f(t_{k+1})\cdots f(t_{l})\,\mathrm{d}t_{l}\cdots\,\mathrm{d}t_{k+1}$
@@
Using the simple combinatorial identity
@@
$\displaystyle\,=e^{-k\varphi(t)}\frac{(-1)^{l-k}}{(l-k)!}\Big{(}k(k+1)\cdots l\Big{)}\int_{0}^{x}(y-x)^{l-k}(1-y)^{-l-1}\,\mathrm{d}y\Big{|}_{x=1-e^{-\varphi(t)}}.$
@@
where $Y$ is some random variable with the Beta distribution $B(k,l-k+1)$. By the sub-Gaussian tail estimate of the Beta distribution in [39, Theorem 2.1], which reads
@@ 3. Regularity Estimates
After presenting the proof of our main results, we are now turning to providing the key logarithmic growth estimates of the limit density, which are essentially applied in the previous discussions. In this section, we shall first collect some basic regularity estimates of the limit density $g$ as ingredients of our logarithmic estimates.
First we borrow some long time asymptotic decay regularity estimates of the vorticity function $\omega$ from the classical work of Kato [32], since $\omega$ solves the vorticity formulation of the 2D Navier-Stokes equation with an $L^{1}$ initial data, as pointed out before. This will give us enough knowledge about the coefficient of the drift term for the limit equation (1.7).
@@
Assume that $g(t,m,x)$ solves the nonlinear Fokker-Planck limit equation (1.7) with initial data $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}(\mathbb{D})$, then the vorticity function $\omega(t,x)$, which is given by the expectation of $g(t,m,x)$ with respect to the circulation variable $m$:
$\omega(t,x)=\int_{\mathbb{R}}mg(t,m,x)\,\mathrm{d}m,$
satisfies the asymptotic decay estimates:
$\|\nabla^{k}K\ast\omega(t,\cdot)\|_{L^{\infty}}\leq M_{k}(1+t)^{-\frac{1+k}{2}}$
for some constants $M_{k}$ for any $0\leq k\leq 2$.
@@
The long time case is directly from [32, Theorem II] and the short time case is proved by the parabolic maximum principle, see for instance [16, Lemma 2.2]. ∎
Next we shall apply the classical Carlen-Loss argument [10] to provide some Gaussian upper bounds and $L^{p}$ norm bounds with respect to the variable $x$ on the limit density $g(t,m,x)$, especially giving the explicit decay rate of those quantities in time $t$, which are optimal if one checks with the easiest heat kernel case. This result is due to the fact that the drift term $K\ast\omega$ is divergence-free and satisfies the long time asymptotic decay rate as in the above Lemma 3.1.
@@
Assume that $g(t,m,x)$ solves the nonlinear Fokker-Planck limit equation (1.7) with initial data $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}(\mathbb{D})$. Assume as in our main results that $g_{0}$ satisfies the Gaussian upper bound
@@
Note that the drift coefficient $K\ast\omega$ in (1.7) is divergence-free, with
$\|K\ast\omega\|_{L^{\infty}}\lesssim t^{-\frac{1}{2}}$
by Lemma 3.1. Hence we can apply the optimal spatial decay result [10, Theorem 3] with $\beta=1/2$ for instance, say
$g(t,m,x)\leq\frac{C}{t}\int_{{\mathbb{R}}^{2}}\exp\Big{(}-\frac{|x-y|^{2}}{8t}\Big{)}g_{0}(m,y)\,\mathrm{d}y.$
@@
$\displaystyle g(t,m,x)\leq\,$ $\displaystyle\frac{C}{t}\int_{{\mathbb{R}}^{2}}\exp\Big{(}-\frac{|x-y|^{2}}{8t}-\frac{|y|^{2}}{C_{0}}\Big{)}\,\mathrm{d}y$
@@
The long time $L^{p}$ norm control is obtained by the optimal smoothing result [10, Theorem 1], while the short time $L^{p}$ norm control follows from the interpolation inequality and the fact that
$\|g(t,\cdot)\|_{L^{1}}\equiv 1,\quad\|g(t,\cdot)\|_{L^{\infty}}\leq\|g_{0}\|_{L^{\infty}}\leq C$
@@
Then we apply the estimates of the fundamental solution to diffusion processes of generalized divergence form from Osada [44] to provide some Gaussian lower bound on the limit density.
@@
Assume that $g(t,m,x)$ solves the nonlinear Fokker-Planck limit equation (1.7) with initial data $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}(\mathbb{D})$. Assume as in our main results that $g_{0}$ satisfies the logarithmic gradient estimate
@@
It is straightforward to check that the limit equation (1.7) belongs to the generalized divergence form in [44], since the diffusion is uniformly elliptic and the drift is divergence-free and uniformly bounded. Then according to [44], we have the Gaussian lower bound for the fundamental solution, denoted by $p(s,x,t,y)$:
$p(s,x,t,y)\geq\frac{1}{C(t-s)}\exp\Big{(}-C\frac{|x-y|^{2}}{t-s}\Big{)}.$
The logarithmic gradient estimate assumption, together with the mean value inequality, yields that $g_{0}$ has Gaussian lower bound
$g_{0}(m,x)\geq\lambda^{-1}\exp(-\lambda|x|^{2})$
with some $\lambda>0$. We conclude that
$\displaystyle g(t,m,x)=$ $\displaystyle\,\int_{\mathbb{D}}p(0,x,t,y)g_{0}(m,y)\,\mathrm{d}y$
@@
[…]
Remark 3.1.
It is straightforward to check that we only need to impose the Gaussian lower bound assumption on the initial data to establish Lemma 3.3. The logarithmic gradient estimate assumption is stated to keep consistency with the main result.
@@
Assume that $g(t,m,x)$ solves the nonlinear Fokker-Planck limit equation (1.7) with initial data $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}(\mathbb{D})$. Then we have for any $1\leq p\leq\infty$,
$\|\nabla g(t,m,\cdot)\|_{L^{p}}\leq Ce^{Ct},$
$\|\nabla^{2}g(t,m,\cdot)\|_{L^{p}}\leq C(1+t)e^{Ct}$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\nabla g\|_{L^{\infty}}\leq C\|K\ast\nabla\omega\|_{L^{\infty}}\|\nabla g\|_{L^{\infty}}\leq C\|\nabla g\|_{L^{\infty}},$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\nabla g\|_{L^{1}}\leq C\|K\ast\nabla\omega\|_{L^{\infty}}\|\nabla g\|_{L^{1}}\leq C\|\nabla g\|_{L^{1}}.$
@@
By the parabolic maximum principle, the evolution of $\|\nabla^{2}g\|_{L^{\infty}}$ is given by
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\nabla^{2}g\|_{L^{\infty}}\leq C\|K\ast\nabla\omega\|_{L^{\infty}}\|\nabla^{2}g\|_{L^{\infty}}+C\|K\ast\nabla^{2}\omega\|_{L^{\infty}}\|\nabla g\|_{L^{\infty}}\leq C\|\nabla^{2}g\|_{L^{\infty}}+Ce^{Ct},$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\nabla^{2}g\|_{L^{1}}\leq C\|K\ast\nabla\omega\|_{L^{\infty}}\|\nabla^{2}g\|_{L^{1}}+C\|K\ast\nabla^{2}\omega\|_{L^{\infty}}\|\nabla g\|_{L^{1}}\leq C\|\nabla^{2}g\|_{L^{1}}+Ce^{Ct}.$
@@
4. Sharp Logarithmic Growth Estimates
In this section, we focus on the key logarithmic growth estimates stated before. As observed in many previous works for instance [16, 11, 48], the major difficulty one will face when working on the whole space instead of the compact domain like the torus, is the spatial growth estimates for the derivatives of the logarithm of the limit density. Roughly speaking, when controlling the error term appearing in the time evolution of the relative entropy, we need to set up some growth estimates for quantities like $\nabla\log\bar{\rho}$ and $\nabla^{2}\log\bar{\rho}$, where $\bar{\rho}$ generally represents the limit density. Such estimates are also crucial when considering the stability of the limit density in the sense of relative entropy, which goes beyond the range of particle approximation.
For the torus case, as in [30, 22, 52], one may assume that the initial density has some positive lower bound $\lambda>0$. By the parabolic maximum principle, the limit density keeps the lower bound for any time, giving the trivial estimate
@@ 4. Sharp Logarithmic Growth Estimates
This assumption even leads to a time-uniform LSI for the limit density using the classical Bakry-Émery arguments, which plays a crucial role in [22] in proving the uniform-in-time propagation of chaos.
But for our whole space case, one can never find such positive lower bound for the density. Therefore, the authors introduced some ideas in [16] to set up the logarithmic gradient and Hessian estimates and started the program of extending the quantitative propagation of chaos results to the whole Euclidean space. These results were obtained using the famous Li-Yau estimate [36] and Hamilton estimate [23, 24], and the Bernstein method with the Grigor’yan maximum principle [21, 20, 37]. A simple parabolic maximum principle method is available for the Landau equation with Maxwellian molecules [11]. In this section, we shall set up the logarithmic estimates for the equation (1.7), and we have simplified our method using only the Grigor’yan parabolic maximum principle, thanks to the Gaussian upper and lower bound results in Lemma 3.2 and Lemma 3.3. We have also optimized the time dependence of the coefficients in the estimate to be sharp in time, improving the bounds in [16].
Proposition 4.1 (Sharp Logarithmic Growth Estimates).
Assume that $g(t,m,x)$ solves the nonlinear Fokker-Planck limit equation (1.7) with initial data $g_{0}\in W_{x}^{2,1}\cap W_{x}^{2,\infty}(\mathbb{D})$. Assume as in our main results that $g_{0}$ satisfies the logarithmic growth estimates
$|\nabla\log g_{0}(m,x)|^{2}\lesssim 1+|x|^{2},$
$|\nabla^{2}\log g_{0}(m,x)|\lesssim 1+|x|^{2},$
@@
$|\nabla\log g(t,m,x)|^{2}\leq\frac{B_{1}}{1+t}(B_{2}-\log(1+t)-\log g(t,m,x)),$ (4.1)
$|\nabla^{2}\log g(t,m,x)|\leq\frac{B_{3}}{1+t}(B_{4}-\log(1+t)-\log g(t,m,x)),$ (4.2)
for some large constants $B_{1},B_{2},B_{3},B_{4}$. Combining with the Gaussian lower bound in Lemma 3.3, we deduce that
$|\nabla\log g(t,m,x)|^{2}\lesssim\frac{1}{1+t}+\frac{|x|^{2}}{(1+t)^{2}},$ (4.3)
$|\nabla^{2}\log g(t,m,x)|\lesssim\frac{1}{1+t}+\frac{|x|^{2}}{(1+t)^{2}}.$ (4.4)
These estimates are now sharp in time, since they coincide with the estimates of the heat kernel and the Oseen vortex.
The basic idea of the proof of is almost the same as in [16], namely applying the Bernstein’s method by creating some auxiliary functions and propagating them along the parabolic equation (1.7). Using the specific structure of the coefficients of the equation, we are able to apply the Grigor’yan parabolic maximum principle [21, 20] to deduce the non-positivity of those auxiliary functions, leading to our desired estimates. However, in order to obtain the sharp time dependence of the estimate, we have modified the proof using some new techniques motivated by [23].